OpenStudy (anonymous):

To create a breakfast beverage, kiwi juice in two concentrations, 14% and 40%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 60 gallons of 14% juice is used, how many gallons of the 40% juice must be used to obtain a 30% kiwi juice solution?

OpenStudy (anonymous):

14% of 60 gallons + 40% of X gallons= 30% of (60+X) gallons or 14% of 60 gallons + 40% of X gallons= 30% of 60 gallons +30% of X gallons or (40-30)% of X gallons= (30-14)% of 60 gallons or 10% of X=16% of 60 10% of X= 9.6 gallons or X=96 gallons

OpenStudy (anonymous):

there is two anwsers when i compute it..........

OpenStudy (anonymous):

Show me the steps by which u worked it out

OpenStudy (anonymous):

i do not understand how you sent up the problem. i need step by step .....

OpenStudy (anonymous):

ok when u take 60 gallons of 14% juice, the amount of juice you actually have is 14% of 60, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. so now when u take X gallons of 40% u have 40% of X..so the total amount of juice with u is 40% of X + 14% of 60

OpenStudy (anonymous):

14% of 60 is 8.4 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

now this amount of juice should be equal to 30% of the total liquid that you have.is this part clear?

OpenStudy (anonymous):

how do you find the equal to 30%?

OpenStudy (anonymous):

ok so the total amount of liquid (not just juice but total liquid) with u is now 60+X gallons..so whatever amount of juice you have must be equal to 30% of this 60+X amount

OpenStudy (anonymous):

can you just give me the anwser please :/

OpenStudy (anonymous):

answer is 96 gallons of 40% juice